OpenStudy (anonymous):

prove that (tan N+cot N)^2 = sec^2 N+csc^2N

OpenStudy (anonymous):

(tan N+cot N)^2 =tan^2N+2cotNtanN+cot^2N =tan^2N+2+cot^2N ==tan^2N+1+1+cot^2N =sec^2N+csc^2N

OpenStudy (anonymous):

(( sin N/cos N + cosN/sinN) )^2 ( sin^2N +cos^2N)^2)/sinN cosN 1/sin^2Ncos^2n cosec^2N sec^2N